# Non-Gravitational Forces in Planetary Systems

### v1

@@ Abstract
Non-gravitational forces play surprising and, sometimes, centrally important roles in shaping the motions and properties of small planetary bodies. In the solar system, the morphologies of comets, the delivery of meteorites and the shapes and dynamics of asteroids are all affected by non-gravitational forces. Around other stars, non-gravitational forces affect the lifetimes of particles and their rates of radial transport within circumstellar disks. Unlike the gravitational force, which is a simple function of the well known separations and masses of bodies, the non-gravitational forces are frequently functions of poorly known or even unmeasurable physical properties. Here, we present order-of-magnitude descriptions of non-gravitational forces, with examples of their application.
@@ 5.1 Tidal Dissipation
The work done by a force $\delta F$ applied over a distance $\delta a_{p}$ is $\delta W\sim\delta F\delta a_{p}$. Substituting gives
@@ 5.2 Internal Dissipation
Assuming $\mu=10^{11}$ N m-2, $Q$ = 100, $\rho$ = 1000 kg m-3 and a middle value $A=30$, we rewrite Equation 22 for the damping time in Myr as
@@
with $a$ in km and rotation period $P_{r}$ in hours. A 1 km body with an excited rotation at $P_{r}$ = 3 hours would damp on timescale $\tau_{damp}\sim$ 300 Myr, by this relation. The 0.1 km scale interstellar object 1I/’Oumuamua, with period $\sim$8 hour, has $\tau_{damp}$ longer than the age of the universe, consistent with photometric evidence that it might be in an excited rotational state (Drahus et al., 2018). Larger asteroids whose rotation should be damped by Equation 23 can nevertheless be excited by impact, radiation torques or gravitational torques from close approaches to planets. The nuclei of some comets also exhibit excited rotation, notably the 15 km long nucleus of 1P/Halley (Samarasinha & A’Hearn, 1991). In the comets, rotational excitation is a natural product of strong sublimation torques. Again, we emphasize that the material properties $\mu$, $Q$ and $A$ (Equation 22) are extremely poorly known (c.f., section 5.3, Figure 4), and damping times very different from Equation 23 are possible.
@@ 5.3 Examples
Figure 4: $\mu Q$ as a function of radius estimated from binary asteroids. The sub-kilometer objects (orange triangles) are near Earth asteroid binaries, for which a median age 107 year is assumed. The $\sim$3 km asteroids in the main belt (yellow circles) are assumed to have median collisional age 109 year while the 100 km objects (green diamonds) are assumed to be 4.5 Gyr, as old as the solar system. Data from Taylor & Margot (2011).
@@
  • Material Properties: Neither $Q$ nor $\mu$ can be calculated from first principles for planetary bodies of interest. Empirically, the quality factor varies widely (e.g., $Q\sim$ 10 for the Earth, $Q\sim$ 100 for the Moon (Goldreich & Soter, 1966)). Values for gas giant planets are uncertain, but probably one or two orders of magnitude lower than $Q\sim 10^{6}$ initially estimated by these authors (c.f., Fuller et al. (2024)). A value $Q\sim$ 100 is often assumed (albeit with little firm evidence) to apply to small solar system bodies. The Young’s moduli of terrestrial rocks span the range from $\mu\sim 10^{9}$ N m-2 for sandstone to $\sim 10^{11}$ N m-2 for basalt to 4×$\times 10^{11}$ N m-2 for diamond. Many small bodies have an aggregate, “rubble pile” structure (Walsh, 2018) that should give them mechanical properties quite different from solid rocks. In particular, the porous and fragmented nature of asteroid rubble piles would seem to point by analogy towards $\mu$ values at the lower end of this terrestrial rock range, in which case $\mu Q\sim 10^{11}$ N m-2 might be expected, but this value is little more than a guess.
@@
    Useful empirical constraints on the product $\mu Q$ can be obtained from the application of Equation 20 to measurements of binary asteroids. For example, if the age of a binary asteroid is known and the other physical properties in the equation are measured or can be estimated, then substitution gives $\mu Q$. Figure 4 shows $\mu Q$ obtained in this way as a function of primary asteroid radius for binaries in three size groups (Taylor & Margot, 2011). The median value for sub-kilometer near-Earth binaries (assuming $\tau_{t}=10^{7}$ yr, the median dynamical lifetime of this population) is $\mu Q=5\times 10^{8}$ N m-2. This is far smaller even than $\mu Q\sim 10^{11}$ N m-2 estimated above. The median value for 3 km scale main-belt binaries (assuming $\tau_{t}=10^{9}$ yr, the approximate collisional lifetime of objects this size) is $\mu Q=7\times 10^{11}$ N m-2. Finally, $\mu Q=1.4\times 10^{11}$ N m-2 for 100 km scale asteroids, computed assuming that they have survived for the 4.5 Gyr age of the solar system. The considerable differences between these estimates may, in part, reflect a size dependence of $\mu Q$ (c.f. Nimmo & Matsuyama (2019)), but also testify to an element of guesswork in assigning values to some of the parameters in Equation 20. For example, it is difficult to assign more than a statistical age (e.g., based on the assumed collisional or dynamical lifetime) to any given binary object.
@@
  • Rotational Dissipation: By Equation 20, the tidal evolution timescale grows particularly strongly (as $r^{6}$) with the separation. In a given system, doubling the separation increases the tidal timescale by a factor 26 = 64, almost two orders of magnitude. This strong distance dependence is evident in real solar system objects. For example, a majority of satellites close to their parent planets rotate synchronously, while the more distant satellites do not. Figure 5 plots the available measurements of the satellites of Saturn, showing the trend for close-in satellites to be rotationally synchronized while the rotations of distant satellites are unrelated to their orbital periods. The 360 km diameter satellite Hyperion stands out. Although Hyperion should be synchronized according to Equation 20, its rotation is instead chaotic (i.e., the period changes irregularly) owing to impulsive torques exerted on its irregular body shape over the course of its eccentric ($e$ = 0.12) orbit (Wisdom et al. (1984), Klavetter (1989a), Klavetter (1989b)).
@@
    The rotation of the Earth’s Moon is obviously synchronized with its orbital motion and measurements show that the Earth’s rotation is slowing (at $\sim 10^{-5}$ s year-1, or a few hours per billion years (Mitchell et al., 2023)) as a result of tidal friction. Inserting $r/a_{p}$ = 60, $m_{p}/m_{s}$ = 81, $\mu Q=10^{11}$ N m-2, $\rho=5\times 10^{3}$ kg m-3 and $a_{p}$ = 6400 km into Equation 20 gives $\tau_{t}\sim$ 0.5 Gyr. Although this is only an order of magnitude estimate, it is clear that tidal interaction with the Moon has slowed the spin of the Earth. Likewise, terrestrial tides on the Moon slowed its rotation on a much shorter ($\sim$10 Myr) timescale, explaining its current synchronous rotation state.
@@ 6 Yarkovsky Force
An isothermal sphere at a finite temperature would radiate photons isotropically, experiencing no net recoil force from the loss of photon momentum. However, real solar system bodies are neither spherical nor isothermal, being hotter on the sun-facing (day) side than on the opposite (night) side. The distribution of slopes, defined by both large and small scale features (e.g., boulders) on the surface, imbues the asteroid with a chirality, so that reflected and radiated photons exert a shape and surface texture dependent torque. Moreover, body rotation, combined with the finite thermal response time of the surface material, carries peak heat from local noon into the afternoon (see Appendix A). As a result of this thermal lag, the afternoon temperatures will always be higher than the corresponding morning temperatures on a rotating body. The chirality and the azimuthal thermal lag asymmetry in the temperature lead to a net force, the so-called diurnal Yarkovsky force. The component of the force parallel to the direction of orbital motion can do work on the orbit, causing it to shrink or expand depending on whether the sense of the rotation is retrograde (i.e., opposite to the sense of orbital motion) or prograde (in the same sense as the orbital motion). See Figure 6.
@@
The above describes the diurnal Yarkovsky effect, which results from axial rotation of the asteroid and is maximized at 0° and 180° obliquity (a geometry assumed, for simplicity, in this discussion and in Figure 6). Given a porous regolith, with diffusivity $\kappa=10^{-9}$ to 10-8 m2 s-1, the thermal skin depth (Equation 2) on an asteroid rotating with period $P$ = 5 hours is $\ell\sim(P\kappa)^{1/2}\sim$ 4 to 13 mm. At non-zero obliquities, the Yarkovsky force is reduced by a factor equal to the cosine of the obliquity. The annual or seasonal Yarkovsky effect has a similar physical origin in thermal lag, but results from motion around the orbit and is maximized instead at 90° obliquity (Figure 7). Whereas the diurnal Yarkovsky effect results from thermal lag on asteroid rotation periods (typically hours), the seasonal effect results from thermal lag on the orbital timescale (typically years). The ratio of seasonal to diurnal timescales is years/hours $\sim$104 and, since the thermal skin depth scales with the square root of the time (Equation 2), the seasonal Yarkovsky effect depends on the thermophysical properties of a surface skin $\sim$102 times deeper than the diurnal effect. Instead of the diurnal Yarkovsky effect being driven by temperature variations in the top $\sim$ 4 to 13 mm of regolith, the seasonal effect is driven by an upper layer $\sim$0.4 to 1.3 m thick. A systematic difference between the seasonal and diurnal effects is that the former always opposes the orbital motion, causing a shrinkage of the orbit, whereas diurnal Yarkovsky can expand or contract it depending on the rotation direction. The relative magnitudes of the diurnal and seasonal Yarkovsky effects depend on many quantities, including the object size, spin vector and the depth dependence of the thermophysical properties of the surface layers (c.f. Appendix A). The diurnal effect typically dominates for bodies with regoliths (i.e., low diffusivity surfaces); we ignore the seasonal effect from further discussion here for simplicity. We also ignore the reflected component of the torque, because the albedos of asteroids are low (C-type albedos are a few percent, while even S-type albedos are only $\sim$20%) and reflected torques are secondary.
@@
Even so, the application to most real asteroids is problematic, because the force depends on many unknown or poorly constrained quantities. These include the magnitude and direction of the spin vector, the body shape, the surface roughness and the thermophysical parameters responsible for the thermal lag (c.f. Equation 1). For all these reasons, the Yarkovsky force cannot in general be calculated for a given body but it can be inferred from careful measurements of the action of the force. The Yarkovsky force has been reviewed by Bottke et al. (2006) and Vokrouhlický et al. (2015). Here we offer an order of magnitude derivation that captures the essence of the process and follow with some examples of its application.
@@
measured in m s-1, and $dr_{H}/dt$ can be positive or negative, depending on the sense of asteroid rotation (c.f. Figure 6). The maximum possible drift rate at 1 au is given by setting $k_{Y}$ = 1, $r_{H}$ = 1 au, with the other parameters as above. For $a$ = 1 km, this gives $d(r_{H})/dt\sim 0.5$ km year-1 or 4×$\times 10^{-3}$ au Myr-1. (We will show below that a more typical value is $k_{Y}\sim$ 0.05, so that $d(r_{H})/dt\sim 2\times 10^{-4}$ au Myr-1 is a better estimate for a 1 km body at 1 au). We conclude that, depending on the actual value of $k_{Y}$, the Yarkovsky force is capable of modifying the orbital properties of kilometer sized asteroids on timescales that are very short compared to the age of the solar system. A practical limit to the influence of Yarkovsky drag is set by collisions with other bodies, which can reset the spin and so change the magnitude and even direction of the radiative torque (Farinella et al. (1998), Wiegert (2015)).
@@ 6.1 Examples
  • Orbital Drift: To obtain an estimate of the Yarkovsky constant, $k_{Y}$, we must use empirical data connecting measurements of the drift rate with the asteroid size, orbit and density. Greenberg et al. (2020) and Fenucci et al. (2024) report drift rate detections for 247 and 348 asteroids, respectively. Figure 8 shows measurements of the magnitude of the radial drift rate as a function of radius for 70 near-Earth asteroids from Fenucci et al. (2024) having semimajor axes near 1 au and formal signal-to-noise ratio $>$10. Lines in the figure show Equation 27 with assumed values $k_{Y}$ = 0.01, 0.1 and 1. The data are reasonably well matched by the expected $1/a$ size dependence of the drift rate (Equation 27) and by $k_{Y}\sim$ 0.02 to 0.13. We take $k_{Y}=0.05$ as a representative value. $k_{Y}\sim$ 0.05 implies a lag angle $\theta=\sin^{-1}(0.05)\sim$ 0.05 radians, or about 3°. It should be remembered that the data are biased towards asteroids with the largest $k_{Y}$, because small values of $k_{Y}$ give small and difficult to measure (low signal to noise ratio) drift rates.
    Figure 8: Absolute value of the radial drift rate plotted as a function of asteroid radius in meters. Lines show the model drift rate calculated from Equation 27 with $\rho=10^{3}$ kg m-3, $r_{H}$ = 1 au, $k_{Y}$ = 0.01 (red dashed line), 0.1 (solid black line) and 1 (dashed blue line). Evidently, $k_{Y}\sim$ 0.05 matches the bulk of the measurements. Data are from the compilation by Fenucci et al. (2024). Error bars are not plotted for clarity of presentation.
    Given a random distribution of asteroid spin vectors, the above model predicts that the ratio of positive (corresponding to prograde rotators) to negative (retrograde rotator) values of $dr_{H}/dt$ should be close to unity. Instead, 70% of the high quality determinations have $dr_{H}/dt<$ 0 (Fenucci et al., 2024), giving a ratio negative drift/positive drift = 2.3 (Figure 9). Greenberg et al. (2020) also found this negative/positive ratio to be $\sim$2.3, indicating an excess of retrograde rotators. This inference matches measurements of the spins of near-Earth asteroids, which show a preponderance of retrograde rotators (retrograde/prograde ratio $\sim$2) (La Spina et al., 2004). The bias in favor of retrograde rotation in the near-Earth population is thought to occur because only inward drifting main belt asteroids can reach the $\nu_{6}$ secular resonance responsible for deflecting them into near-Earth space.
@@
  • Dispersal of Asteroid Families: Shattering collisions between asteroids produce families of objects having initially similar orbital elements. However, family asteroids experience a size dependent drift from their initial orbital semimajor axes under the action of the Yarkovsky force. By Equation 27, the semimajor axis drift in time $\Delta t$ is $\Delta r_{H}\propto\Delta t/a$, and $\Delta r_{H}$ can be positive or negative depending on the sense of rotation of the asteroid. Small asteroids drift the farthest, all else being equal, giving rise to a V-shaped distribution in a plot of semimajor axis vs. $1/a$. This effect is observed in the main asteroid belt, where it has been used to identify families and estimate their ages. An example is shown in Figure 10 which compares data for the Erigone family with Equation 27 (we assumed $k_{Y}$ = 0.05, $\rho$ = 2000 kg m-3 and initial semimajor axis $r_{H}$ = 2.37 au). There is some ambiguity in defining the edges of the distribution, but spreading ages near 70 Myr to 100 Myr are plausible, while much larger and smaller ages are not. Bolin et al. (2018) estimated an age of 90 Myr. The method is not perfect. The initial orbits may not have exactly the same semimajor axis, the effect of asteroid obliquity is not included (Equation 27 gives the maximum value, for obliquity = 0°) and the individual obliquities are in any case not known, the sizes of asteroids are approximate, and external perturbations (e.g., from planets) can act to spread the orbits (especially for the older clusters). Nevertheless, Yarkovsky spreading gives an impressive explanation for the V-shaped diagram and a useful measure of the family age.
@@
  • Post-Main Sequence Evolution of the Sun: In $\lesssim$5 Gyr the Sun will exhaust its core hydrogen and bloat into a red giant, with a luminosity $\sim 10^{3.5}$ times its present value (Vassiliadis & Wood, 1993). When this occurs, the Yarkovsky force will likewise jump by a factor $\sim$103.5 causing accelerated orbital drift to compete with many other processes to destabilize small orbiting bodies of the solar system. Destabilized planetary systems are likely implicated in feeding debris to the central white dwarf remaining after the red giant phase (Veras et al., 2019).
@@
  • Binary Yarkovsky: The Yarkovsky force can affect a binary in two ways. First, in a binary with small separation, thermal radiation from each component can heat the other, resulting in a kind of radiation drag distinct from the Yarkovsky drag caused by direct sunlight. Second, if the orbital inclination is also small, the two components will pass through each other’s shadow in each orbit. Since the thermal response of the surface material is lagged (see Appendix A), the entry into and exit from the shadow introduces an additional asymmetry into the radiative force that results in another torque. These subtle effects, the former somewhat unfortunately named “planetary Yarkovsky” and the latter “Yarkovsky-Schach” drag) may rival tidal dissipation (Section 5.1) and BYORP (Section 8.2) in the orbital evolution of some close binary asteroids (Zhou et al., 2024).

### v2

@@ Abstract
Non-gravitational forces play surprising and, sometimes, centrally important roles in shaping the motions and properties of small planetary bodies. In the solar system, the morphologies of comets, the delivery of meteorites and the shapes and dynamics of asteroids and binaries are all affected by non-gravitational forces. In exoplanetary systems and debris disks, non-gravitational forces affect the lifetimes of circumstellar particles and feed refractory debris to the photospheres of the central stars. Unlike the gravitational force, which is a simple function of the well known separations and masses of bodies, the non-gravitational forces are frequently functions of poorly known or even unmeasurable physical properties. Here, we present order-of-magnitude descriptions of non-gravitational forces, with examples of their application.
@@ 5.1 Tidal Dissipation
The work done by a force $\delta F$ applied over a distance $\delta a_{p}$ is $\delta W\sim\delta F\delta a_{p}$. Substituting from Equations 16 and 17 gives
@@ 5.2 Internal Dissipation
Assuming $\mu=10^{11}$ N m-2, $Q$ = 100, $\rho$ = 1000 kg m-3 and a geometric middle value $A=30$, we rewrite Equation 22 for the damping time in Myr as
@@
with $a$ in km and rotation period $P_{r}$ in hours. A 1 km body with an excited rotation at $P_{r}$ = 3 hours would damp on timescale $\tau_{damp}\sim$ 300 Myr, by this relation. The 0.1 km scale interstellar object 1I/’Oumuamua, with period $\sim$8 hour, has $\tau_{damp}$ longer than the age of the universe, consistent with photometric evidence that it might be in an excited rotational state (Drahus et al., 2018). Larger asteroids whose rotation should be damped according to Equation 23 can nevertheless be excited by impact, radiation torques or gravitational torques from close approaches to planets. The nuclei of some comets also exhibit excited rotation, notably the 15 km long nucleus of 1P/Halley (Samarasinha & A’Hearn, 1991). In the comets, rotational excitation is a natural product of strong sublimation torques. Again, we emphasize that the material properties $\mu$, $Q$ and $A$ (Equation 22) are extremely poorly known (c.f., section 5.3, Figure 4), and damping times very different from those given by Equation 23 are possible.
@@ 5.3 Examples
Figure 4: $\mu Q$ as a function of radius estimated from binary asteroids. The sub-kilometer objects (orange triangles) are near Earth asteroid binaries, for which a median age 107 year is assumed. The $\sim$3 km asteroids in the main belt (yellow circles) are assumed to have median collisional age 109 year while the 100 km objects (green diamonds) are assumed to be 4.5 Gyr, as old as the solar system. Note the enormous scatter in $\mu Q$. Data from Taylor & Margot (2011).
@@
  • Material Properties: Neither $Q$ nor $\mu$ can be calculated from first principles for planetary bodies of interest. Empirically, the quality factor varies widely (e.g., $Q\sim$ 10 for the Earth, $Q\sim$ 100 for the Moon (Goldreich & Soter, 1966)). Values for gas giant planets are uncertain, but probably one or two orders of magnitude lower than $Q\sim 10^{6}$ initially estimated by these authors (c.f., Fuller et al. (2024)). A value $Q\sim$ 100 is often assumed (albeit with little firm evidence) to apply to small solar system bodies. Separately, the Young’s moduli of terrestrial rocks span the range from $\mu\sim 10^{9}$ N m-2 for sandstone and $\sim 10^{11}$ N m-2 for basalt to 4×$\times 10^{11}$ N m-2 for diamond. Many small bodies have an aggregate, “rubble pile” structure (Walsh, 2018) that should give them mechanical properties quite different from solid rocks. In particular, the porous and fragmented nature of asteroid rubble piles would seem to point by analogy towards $\mu$ values at the lower end of this terrestrial rock range, in which case $\mu Q\sim 10^{11}$ N m-2 might be expected. This value remains little more than a guess, and evidence from asteroid binaries suggests that even smaller $\mu Q$ values may prevail, as we next describe.
@@
    Useful empirical constraints on the product $\mu Q$ can be obtained from the application of Equation 20 to measurements of binary asteroids. For example, if the age of a binary asteroid is known and the other physical properties (separation, period, component sizes, and density) in the equation are measured or can be estimated, then substitution gives $\mu Q$. Figure 4 shows $\mu Q$ obtained in this way as a function of primary asteroid radius for binaries in three size groups (Taylor & Margot, 2011). The median value for sub-kilometer near-Earth binaries (assuming $\tau_{t}=10^{7}$ yr, the median dynamical lifetime of this population) is $\mu Q=5\times 10^{8}$ N m-2. This is far smaller even than $\mu Q\sim 10^{11}$ N m-2 estimated above. The median value for 3 km scale main-belt binaries (assuming $\tau_{t}=10^{9}$ yr, the approximate collisional lifetime of objects this size) is $\mu Q=7\times 10^{11}$ N m-2. Finally, $\mu Q=1.4\times 10^{11}$ N m-2 for 100 km scale asteroids, computed assuming that they have survived for the 4.5 Gyr age of the solar system. The considerable differences between these estimates may, in part, reflect a real size dependence of $\mu Q$ (c.f. Nimmo & Matsuyama (2019)), but they also testify to an element of guesswork in assigning values to some of the parameters in Equation 20. For example, it is difficult to assign more than a statistical age (e.g., based on the assumed collisional or dynamical lifetime) to any given binary object.
@@
  • Rotational Dissipation: By Equation 20, the tidal evolution timescale grows particularly strongly (as $r^{6}$) with the separation. In a given system, doubling the separation increases the tidal timescale by a factor 26 = 64, almost two orders of magnitude. This strong distance dependence is evident in real solar system objects. For example, a majority of satellites close to their parent planets rotate synchronously, while the more distant satellites do not. Figure 5 plots the available measurements of the satellites of Saturn, showing the trend for close-in satellites to be rotationally synchronized while the rotations of distant satellites are unrelated to their orbital periods. The 360 km diameter satellite Hyperion stands out. Although Hyperion should be synchronized according to Equation 20, its rotation is instead chaotic (i.e., the period changes irregularly) owing to impulsive torques exerted on its irregular body shape over the course of its eccentric ($e$ = 0.12) orbit, as predicted by Wisdom et al. (1984) and established by Klavetter (1989a) and Klavetter (1989b).
@@
    The rotation of the Earth’s Moon is obviously synchronized with its orbital motion and measurements show that the Earth’s rotation is currently slowing (at $\sim 10^{-5}$ s year-1, or $\sim$3 hours per billion years (Mitchell et al., 2023)) as a result of tidal friction. Inserting $r/a_{p}$ = 60, $m_{p}/m_{s}$ = 81, $\mu Q=10^{11}$ N m-2, $\rho=5\times 10^{3}$ kg m-3 and $a_{p}$ = 6400 km into Equation 20 gives $\tau_{t}\sim$ 0.5 Gyr. Although this is only accurate to an order of magnitude, it is clear that tidal interaction with the Moon is responsible for slowing the spin of the Earth. Likewise, terrestrial tides on the Moon slowed its rotation on a much shorter ($\sim$10 Myr) timescale, explaining its current synchronous rotation state.
@@ 6 Yarkovsky Force
An isothermal sphere at a finite temperature would radiate photons isotropically, experiencing no net recoil force from the loss of photon momentum. However, real solar system bodies are neither spherical nor isothermal, being hotter on the sun-facing (day) side than on the opposite (night) side. The distribution of slopes, defined by both large and small scale features (e.g., boulders) on the surface, imbues the asteroid with a chirality, so that reflected and radiated photons exert a shape and surface texture dependent torque. Moreover, body rotation, combined with the finite thermal response time of the surface material, carries peak heat from local noon into the afternoon, giving a “thermal lag” (see Appendix A). As a result of this thermal lag, the afternoon temperatures will always be higher than the corresponding morning temperatures on a rotating body. The chirality and the azimuthal thermal lag asymmetry in the temperature lead to a net force, the so-called diurnal Yarkovsky force. The component of the force parallel to the direction of orbital motion can do work on the orbit, causing it to shrink or expand depending on whether the sense of the rotation is retrograde (i.e., opposite to the sense of orbital motion) or prograde (in the same sense as the orbital motion). See Figure 6.
@@
The above describes the diurnal Yarkovsky effect, which results from axial rotation of the asteroid and is maximized at 0° and 180° obliquity (a geometry assumed, for simplicity, in this discussion and in Figure 6). For example, given a porous regolith with diffusivity $\kappa=10^{-9}$ to 10-8 m2 s-1, the thermal skin depth (Equation 2) on an asteroid rotating with period $P$ = 5 hours is $\ell\sim(P\kappa)^{1/2}\sim$ 4 to 13 mm. At non-zero obliquities, the Yarkovsky force is reduced by a factor equal to the cosine of the obliquity.
The annual or seasonal Yarkovsky effect has a similar physical origin in thermal lag, but results from motion around the orbit and is maximized instead at 90° obliquity (Figure 7). One important, systematic difference between the seasonal and diurnal effects is that the former always opposes the orbital motion, causing a shrinkage of the orbit, whereas diurnal Yarkovsky can expand or contract it depending on the rotation direction. A second difference is that, whereas the diurnal Yarkovsky effect results from thermal lag on asteroid rotation periods (typically hours), the seasonal effect results from thermal lag on the orbital timescale (typically years). The ratio of seasonal to diurnal timescales is years/hours $\sim$104 and, since the thermal skin depth scales with the square root of the time (Equation 2), the seasonal Yarkovsky effect depends on the thermophysical properties of a surface skin $\sim$102 times deeper than the diurnal Yarkovsky. Instead of the diurnal Yarkovsky effect being driven by temperature variations in the top $\sim$ 4 to 13 mm of regolith, the seasonal effect is driven by an upper layer $\sim$0.4 to 1.3 m thick, all else being equal. The relative magnitudes of the diurnal and seasonal Yarkovsky effects depend on many quantities, including the object size, spin vector and the depth dependence of the thermophysical properties of the surface layers (c.f. Appendix A). The diurnal effect typically dominates for bodies with regoliths (i.e., low diffusivity surfaces); we ignore the seasonal effect from further discussion here for simplicity. We also ignore the reflected component of the torque, because the albedos of asteroids are low (C-type albedos are a few percent, while even S-type albedos are only $\sim$20%) and reflected torques are secondary.
@@
Even so, the application to most real asteroids is problematic, because the force depends on many unknown or poorly constrained quantities. These include the magnitude and direction of the spin vector, the body shape, the surface roughness and the thermophysical parameters responsible for the thermal lag (c.f. Equation 1). For all these reasons, the Yarkovsky force cannot in general be calculated for a given body. Instead, it can be inferred from careful measurements of the action of the force. The Yarkovsky force has been reviewed by Bottke et al. (2006) and Vokrouhlický et al. (2015). Here we offer an order of magnitude derivation that captures the essence of the process and we follow with some examples of its application.
@@
measured in m s-1, and $dr_{H}/dt$ can be positive or negative, depending on the sense of asteroid rotation (c.f. Figure 6). The maximum possible drift rate at 1 au is given by setting $k_{Y}$ = 1, $r_{H}$ = 1 au, with the other parameters as above. For $a$ = 1 km, this gives $d(r_{H})/dt\sim 0.5$ km year-1 or 4×$\times 10^{-3}$ au Myr-1. (We will show below that a more typical value is $k_{Y}\sim$ 0.05, so that $d(r_{H})/dt\sim 2\times 10^{-4}$ au Myr-1 is a better estimate for a 1 km body at 1 au). The asteroid belt is dynamically structured (e.g., at resonance locations) on scales $<$0.05 au, which can be crossed by kilometer-sized asteroids on timescales $<$250 Myr. We conclude that, depending on the size of the asteroid and the actual value of $k_{Y}$, the Yarkovsky force is capable of modifying the orbital properties on timescales that are very short compared to the age of the solar system. A practical limit to the influence of Yarkovsky drag is set by collisions with other bodies, which can reset the spin and so change the magnitude and even direction of the radiative torque (Farinella et al. (1998), Wiegert (2015)).
@@ 6.1 Examples
  • Orbital Drift: To obtain an estimate of the Yarkovsky constant, $k_{Y}$, we must use empirical data connecting measurements of the drift rate with the asteroid size, orbit and density. Greenberg et al. (2020) and Fenucci et al. (2024) report drift rate detections for 247 and 348 asteroids, respectively. Figure 8 shows measurements of the magnitude of the radial drift rate as a function of radius for 58 near-Earth asteroids from Fenucci et al. (2024), having semimajor axes near 1 au and formal signal-to-noise ratio $>$10. Lines in the figure show Equation 27 with assumed values $k_{Y}$ = 0.01, 0.1 and 1. The data are reasonably well matched by the expected $1/a$ size dependence of the drift rate (Equation 27) and by $k_{Y}\sim$ 0.02 to 0.13. We take $k_{Y}=0.05$ as a representative value. $k_{Y}\sim$ 0.05 implies a lag angle $\theta=\sin^{-1}(0.05)\sim$ 0.05 radians, or about 3°. It should be remembered that the data are biased towards asteroids with the largest $k_{Y}$, because small values of $k_{Y}$ give small and difficult to measure (low signal to noise ratio) drift rates.
    Figure 8: Absolute value of the radial drift rate plotted as a function of asteroid radius in meters. Lines show the model drift rate calculated from Equation 27 with $\rho=10^{3}$ kg m-3, $r_{H}$ = 1 au, $k_{Y}$ = 0.01 (red dashed line), 0.1 (solid black line) and 1 (dashed blue line). Evidently, the data are broadly consistent with $k_{Y}\sim$ 0.05. Data are from the compilation by Fenucci et al. (2024).
    Given a random distribution of asteroid spin vectors, the above model predicts that the ratio of positive (corresponding to prograde rotators) to negative (retrograde rotator) values of $dr_{H}/dt$ should be close to unity. Instead, in both the studies by Greenberg et al. (2020) and by Fenucci et al. (2024), fully 70% of the high quality determinations have $dr_{H}/dt<$ 0, giving a ratio negative drift/positive drift = 2.3, and indicating an excess of retrograde rotators. (Figure 9). This inference matches measurements of the spins of near-Earth asteroids, which show a preponderance of retrograde rotators (retrograde/prograde ratio $\sim$2) (La Spina et al., 2004). The bias in favor of retrograde rotation in the near-Earth population is thought to occur because inward drifting main belt asteroids are more easily able to reach the $\nu_{6}$ secular resonance responsible for their deflection into near-Earth space.
@@
  • Dispersal of Asteroid Families: Shattering collisions between asteroids produce families of objects having initially similar orbital elements. However, family asteroids experience a size dependent drift from their initial orbital semimajor axes under the action of the Yarkovsky force. By Equation 27, the semimajor axis drift in time $\Delta t$ is $\Delta r_{H}\propto\Delta t/a$, and $\Delta r_{H}$ can be positive or negative depending on the sense of rotation of the asteroid. Small asteroids drift the farthest, all else being equal, giving rise to a V-shaped distribution in a plot of semimajor axis vs. $1/a$. This effect is observed in the main asteroid belt, where it has been used to identify families and estimate their ages. An example is shown in Figure 10 which compares data for the Erigone family with Equation 27 (we assumed $k_{Y}$ = 0.05, $\rho$ = 2000 kg m-3 and initial semimajor axis $r_{H}$ = 2.37 au). There is some ambiguity in defining the edges of the distribution, but spreading ages near 70 Myr to 100 Myr are plausible, while much larger and smaller ages are not. Bolin et al. (2018) estimated an age of 90 Myr. The method is not perfect. The initial orbits may not have exactly the same semimajor axis, the effect of asteroid obliquity is not included (Equation 27 gives the maximum value, for obliquity = 0°) and the individual obliquities are in any case not known, the sizes and densities of asteroids are approximate, and external perturbations (e.g., from planets) can act to spread the orbits (especially for the older clusters). Nevertheless, Yarkovsky spreading gives an impressive explanation for the V-shaped diagram and a useful measure of the family age.
@@
  • Post-Main Sequence Evolution of the Sun: In $\lesssim$5 Gyr the Sun will exhaust its core hydrogen and bloat into a red giant, with a luminosity $\sim 10^{3.5}$ times its present value (Vassiliadis & Wood, 1993). When this occurs, the Yarkovsky force will likewise jump by a factor $\sim$103.5 causing accelerated orbital drift to compete with many other processes to destabilize small orbiting bodies of the solar system. Destabilized planetary systems are likely implicated in feeding debris to the metal polluted central white dwarf remaining after the red giant phase (Veras et al., 2019).
@@
  • Binary Yarkovsky: The Yarkovsky force can affect a binary in two ways. First, in a binary with small separation, thermal radiation from each component can heat the other, resulting in a kind of radiation drag distinct from the Yarkovsky drag caused by direct sunlight. Second, if the orbital inclination is also small, the two components will pass through each other’s shadow in each orbit. Since the thermal response of the surface material is lagged (see Appendix A), the entry into and exit from the shadow introduces an additional asymmetry into the radiative force that results in yet another torque. These subtle effects, the former somewhat unfortunately named “planetary Yarkovsky” and the latter “Yarkovsky-Schach” drag, may rival tidal dissipation (Section 5.1) and BYORP (Section 8.2) in the orbital evolution of some close binary asteroids (Zhou et al., 2024).